Find the height of the cylinder whose volume if 1.54 m^3 and diameter of the base is 140 cm.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given: Volume of cylinder = 1.54m^3and Diameter of cylinder = 140 cm
Radius ( r )= d/2 = 140/2 = 70 cm
Volume of cylinder = πr^2h
1.54 = (22/7)×0.7×0.7×h
After simplifying, we get the value of h that is h = (1.54×7)/(22×0.7×0.7)
h =1
Hence, height of the cylinder is 1 m.

Thanks